OpenStudy (loser66):

Please help Prove nth root of 2 is irrational (n>= 2)

OpenStudy (loser66):

We know that \(\sqrt 2\) is irrational \(\sqrt[3] 2\) is irrational as well how about \(\sqrt[4] 2~~,\, \sqrt[5] 2,......\)

OpenStudy (loser66):

@SithsAndGiggles